OpenStudy (anonymous):

If the mean of 6, 4, 7, p and 10 is 8, find the value of p.

OpenStudy (anonymous):

sum of all the terns divided by total no of terms = 9 \[\frac{6+4+7+p+10}{5}= 8\]

OpenStudy (anonymous):

Add them all together and divide by 5 to get the answer, which is 8. \[{6+4+7+p+10 \over 5} = 8\]

OpenStudy (anonymous):

Now find p

OpenStudy (diyadiya):

\[Mean = \frac{Sum \ of \ all \ numbers}{Total \ no \ of \ terms}\]

OpenStudy (anonymous):

Typo:sum of all the terns divided by total no of terms = *8

OpenStudy (anonymous):

p=13

OpenStudy (anonymous):

yes